### v4

# Generalizable, Fast, and Accurate DeepQSPR with fastprop

Abstract
Quantitative Structure-Property Relationship studies (QSPR), often referred to interchangeably as QSAR, seek to establish a mapping between molecular structure and an arbitrary Quantity of Interest (QOI). Historically this was done on a QOI-by-QOI basis with new descriptors being devised by researchers to specifically map to their QOI. A large number of descriptors have been invented, and can be computed using packages like DRAGON (later E-dragon), PaDEL-descriptor (and padelpy), Mordred, CODESSA, and many others. The sheer number of different descriptor packages resulted in the creation of ‘meta-packages’ which served only to aggregate these other calculators, including tools like molfeat, ChemDes, Parameter Client, and AIMSim.
Generalizable descriptor-based modeling was a natural evolution of these meta-packages’ development. Historically QSPR researchers focused almost exclusively on linear methods. Another community of researchers focused on finding nonlinear correlations between molecular structures and a QOI, often using Deep learning (DL). The DL community typically used molecular fingerprints instead of the complex descriptors popular in QSPR community. Recently the DL community has turned to learned representations primarily via message passing graph neural networks. This approach has proved remarkably effective but is not without drawbacks. Learning a representation requires large datasets to avoid over-fitting or even learn at all, loses interpretability since an embedding’s meaning can only be induced retroactively, and needs significant execution time given the complexity of the underlying message passing algorithm. This paper introduces fastprop, a software package and general Deep-QSPR framework that combines a cogent set of molecular descriptors with DL to achieve state-of-the-art performance on datasets ranging from tens to tens of thousands of molecules. fastprop is designed with Research Software Engineering best practices and is free and open source, hosted at github.com/jacksonburns/fastprop.
Scientific Contribution
fastprop is a QSPR framework that achieves state-of-the-art accuracy on datasets of all sizes without sacrificing speed or interpretability. As a software package fastprop emphasizes Research Software Engineering best practices, reproducibility, and ease of use for experts across domains.
Keywords
QSPR
Learned Representations
Deep Learning
Molecular Descriptors
Introduction
Chemists have long sought a method to relate only the connectivity of a molecule to its corresponding molecular properties. The Quantitative Structure-Property Relationship (QSPR) would effectively solve the forward problem of molecular engineering and enable rapid development. Reviews on the topic are numerous and cover an enormous range of scientific subdomains; a comprehensive review of the literature is beyond the scope of this publication, though the work of Muratov and coauthors [1] provides an excellent starting point for further review. An abridged version of the history behind QSPR is presented here to contextualize the approach of fastprop.
Historical Approaches
Early in the history of computing, limited computational power meant that significant human expertise was required to guide QSPR models toward effectiveness. This materialized in the form of bespoke molecular descriptors: the Wiener Index in 1947 [2], Atom-Bond Connectivity indices in 1998 [3], and thousands of others. To this day descriptors are still being developed - the geometric-harmonic-Zagreb degree based descriptors were proposed by Arockiaraj et al. in 2023 [4]. In each case, domain experts devised an algorithm which mapped a molecular structure to some scalar value. This algorithm would take into account features of the molecule which that expert deduced were relevant to the property at hand. This time consuming technique is of course highly effective but the dispersed nature of this chemical knowledge means that these descriptors are spread out throughout many journals and domains with no single source to compute them all.
The range of regression techniques applied to these descriptors has also been limited. As explained by Muratov et. al [1] QSPR uses linear methods (some of which are now called machine learning) almost exclusively. The over-reliance on this category of approaches may be due to priorities; domain experts seek interpretability in their work, especially given that the inputs are physically meaningful descriptors, and linear methods lend themselves well to this approach. Practice may also have been a limitation, since historically training and deploying neural networks required more computer science expertise than linear methods.
All of this is not to say that Deep Learning (DL) has never been applied to QSPR. Applications of DL to QSPR, i.e. DeepQSPR, were attempted throughout this time period but focused on the use of molecular fingerprints rather than descriptors. This may be at least partially attributed to knowledge overlap between deep learning experts and this sub-class of descriptors. Molecular fingerprints are bit vectors which encode the presence or absence of sub-structures in an analogous manner to the “bag of words” featurization strategy common to natural language processing. Experts have bridged this gap to open this subdomain and proved its effectiveness. In Ma and coauthors’ review of DL for QSPR [5], for example, it is claimed that DL with fingerprint descriptors is more effective than with molecular-level descriptors. They also demonstrate that DL outperforms or at least matches classical machine learning methods across a number of ADME-related datasets. The results of this study demonstrate that molecular-level descriptors actually are effective and reaffirm that DL matches or outperforms baselines, in this case linear.
Despite their differences, both classical- and Deep-QSPR shared a lack of generalizability. Beyond the domains of chemistry where many of the descriptors had been originally devised, models were either unsuccessful or more likely simply never evaluated. As interest began to shift toward the prediction of molecular properties which were themselves descriptors (i.e. derived from quantum mechanics simulations) - to which none of the devised molecular descriptors were designed to be correlated - learned representations (LRs) emerged.
Shift to Learned Representations
The exact timing of the transition from fixed descriptors (molecular-level or fingerprints) to LRs is difficult to ascertain [6]. Among the most cited at least is the work of Yang and coworkers in 2019 [7] which laid the groundwork for applying LRs to “Property Prediction” - QSPR by another name. In short, the basic idea is to initialize a molecular graph with only information about its bonds and atoms such as order, aromaticity, atomic number, etc. Then via a Message Passing Neural Network (MPNN) architecture, which is able to aggregate these atom- and bond-level features into a vector in a manner which can be updated, the ‘best’ representation of the molecule is found during training. This method proved highly accurate and achieved the generalizability apparently lacking in descriptor-based modeling. The modern version of the corresponding software package Chemprop (described in [8]) has become a de facto standard for property prediction, partially because of the significant development and maintenance effort supporting that open source software project.
Following the initial success of Chemprop numerous representation learning frameworks have been devised, all of which slightly improve performance. The Communicative-MPNN (CMPNN) framework is a modified version of Chemprop with a different message passing scheme to increase the interactions between node and edge features [9]. Uni-Mol incorporates 3D information and relies extensively on transformers [10]. In a “full circle moment” architectures like the Molecular Hypergraph Neural Network (MHNN) have been devised to learn representations for specific subsets of chemistry, in that case optoelectronic properties [11]. Myriad others exist including GSL-MPP (accounts for intra-dataset molecular similarity) [12], SGGRL (trains three representations simultaneously using different input formats) [13], and MOCO (multiple representations and contrastive pretraining) [14].
Limitations
Despite the continuous incremental performance improvements, this area of research has had serious drawbacks. A thru-theme in these frameworks is the increasing complexity of DL techniques and consequent un-interpretability. This also means that actually using these methods to do research on real-world dataset requires varying amounts of DL expertise, creating a rift between domain experts and these methods. Perhaps the most significant failure is the inability to achieve good predictions on small 111What constitutes a ‘small’ dataset is decidedly not agreed upon by researchers. For the purposes of this study, it will be used to refer to datasets with ~1000 molecules or fewer, which the authors believe better reflects the size of real-world datasets. datasets. This is a long-standing limitation, with the original Chemprop paper stating that linear models are about on par with Chemprop for datasets with fewer than 1000 entries [7].
This limitation is especially challenging because it is a fundamental drawback of the LR approach. Without the use of advanced DL techniques like pre-training or transfer learning, the model is essentially starting from near-zero information every time a model is created. This inherently requires larger datasets to allow the model to effectively ‘re-learn’ the chemical intuition which was built in to descriptor- and fixed fingerprint-based representations.
Efforts are of course underway to address this limitation, though none are broadly successful. One simple but incredibly computationally expensive approach is to use delta learning, which artificially increases dataset size by generating all possible pairs of molecules from the available data (thus squaring the size of the dataset). This was attempted by Nalini et al. [15], who used an unmodified version of Chemprop referred to as ‘DeepDelta’ to predict differences in molecular properties for pairs of molecules. They achieve increased performance over standard LR approaches but lost the ability to train on large datasets due to simple runtime limitations. Other increasingly complex approaches are discussed in the outstanding review by van Tilborg et al. [16].
While iterations on LRs and novel approaches to low-data regimes have been in development, the classical QSPR community has continued their work. A turning point in this domain was the release of mordred, a fast and well-developed package capable of calculating more than 1600 molecular descriptors [17]. Critically this package was fully open source and written in Python, allowing it to readily interoperate with the world-class Python DL software ecosystem that greatly benefitted the LR community. Despite previous claims that molecular descriptors cannot achieve generalizable QSPR in combination with DL, the opposite is shown here with fastprop.
Implementation
At its core the fastprop ‘architecture’ is simply the mordred molecular descriptor calculator 222The original mordred package is no longer maintained. fastprop uses a fork of mordred called mordredcommunity that is maintained by community-contributed patches (see github.com/JacksonBurns/mordred-community). Multiple descriptor calculators from the very thorough review by McGibbon et al. [18] could be used instead, though none are as readily interoperable as mordred. [17] connected to a Feedforward Neural Network (FNN) implemented in PyTorch Lightning [19] (Figure 1) - an existing approach formalized into an easy-to-use, reliable, and correct implementation. fastprop is highly modular for seamless integration into existing workflows and includes end-to-end interfaces for general use. In the latter mode the user simply specifies a set of SMILES [20], a linear textual encoding of molecules, and their corresponding properties. fastprop automatically calculates and caches the corresponding molecular descriptors with mordred, re-scales both the descriptors and the targets appropriately, and then trains an FNN to predict the indicated target properties. By default this FNN is two hidden layers with 1800 neurons each connected by ReLU activation functions, though the configuration can be readily changed via the command line interface or configuration file. fastprop owes its success to the cogent set of descriptors assembled by the developers of mordred, the ease of training FNNs with modern software like PyTorch Lightning, and the careful application of Research Software Engineering best practices that make it as user friendly as the best-maintained alternatives.
This trivially simple idea has been alluded to in previous published work but neither described in detail nor lauded for its generalizability or accuracy. Comesana and coauthors, based on a review of the biofuels property prediction landscape, claimed that methods (DL or otherwise) using large numbers of molecular descriptors were unsuccessful, instead proposing a feature selection method [21]. As a baseline in a study of photovoltaic property prediction, Wu et al. reported using the mordred descriptors in combination with both a Random Forest and an Artificial Neural Network, though in their hands the performance is worse than their bespoke model and no code is available for inspection [22].
Others have also incorporated mordred descriptors into their modeling efforts, though none with a simple FNN as described above. Esaki and coauthors started a QSPR study with mordred descriptors for a dataset of small molecules, but via an enormously complex modeling pipeline (using only linear methods) removed all but 53 [23]. Yalamanchi and coauthors used DL on mordred descriptors as part of a two-headed representation, but their network architecture was sequential hidden layers decreasing in size to only 12 features [24] as opposed to the constant 1800 in fastprop.
The reason fastprop stands out from these studies and contradicts previous reports is for the simple reason that it works. As discussed at length in the Results & Discussion section, this approach matches the performance of leading LR approaches on common benchmark datasets and bespoke QSPR models on small real-world datasets. fastprop also overcomes the limitations of LRs discussed above. Because all inputs to the FNN are physically meaningful molecular descriptors, intermediate representations in the FNN are also physically meaningful and can be directly interpreted. The simplicity of the framework enables domain experts to apply it easily and makes model training dramatically faster than LRs. Most importantly this approach is successful on the smallest of real-world datasets. By starting from such an informed initialization the FNN can be readily trained on datasets with as few as forty training examples (see PAHs).
Example Usage
fastprop is built with ease of use at the forefront of design. To that end, input data is accepted in the immensely popular Comma-Separated Values (CSV) format, editable with all modern spreadsheet editors and completely platform independent. An example specify some properties for benzene is shown below, including its SMILES string:
compound,smiles,log_p,retention_index,boiling_point_c,acentric_factor Benzene,C1=CC=CC=C1,2.04,979,80,0.21
fastprop itself is accessed via the command line interface, with configuration parameters passed as either command line arguments or in an easily editable configuration file:
# pah.yml # generic args output_directory: pah random_seed: 55 problem_type: regression # featurization input_file: pah/arockiaraj_pah_data.csv target_columns: log_p smiles_column: smiles descriptor_set: all # preprocessing clamp_input: True # training number_repeats: 4 number_epochs: 100 batch_size: 64 patience: 15 train_size: 0.8 val_size: 0.1 test_size: 0.1 sampler: random
Training, prediction, and feature importance are then readily accessible via the commands fastprop train, fastprop predict, or fastprop shap, respectively. The fastprop GitHub repository contains a Jupyter notebook runnable from the browser via Google colab which allows users to actually execute the above example, which is also discussed at length in the PAHs section, as well as further details about each configurable option.
Results & Discussion
There are a number of established molecular property prediction benchmarks commonly used in LR studies, especially those standardized by MoleculeNet [25]. Principal among them are QM8 [26] and QM9 [27], often regarded as the standard benchmark for property prediction. These are important benchmarks and QM9 is included for completeness, though the enormous size and rich coverage of chemical space in the QM9 dataset means that nearly all model architectures are highly accurate, including fastprop.
Real world experimental datasets, particularly those common in QSPR studies, often number in the hundreds. To demonstrate the applicability of fastprop to these regimes, many smaller datasets are selected including some from the QSPR literature that are not established benchmarks. These studies relied on more complex and slow modeling techniques (ARA) or the design of a bespoke descriptor (PAHs) and have not yet come to rely on learned representations as a go-to tool. In these data-limited regimes where LRs sometimes struggle, fastprop and its intuition-loaded initialization are highly powerful. To emphasize this point further, the benchmarks are presented in order of dataset size, descending.
Two additional benchmarks showing the limitations of fastprop are included after the main group of benchmarks: Fubrain and QuantumScents. The former demonstrates how fastprop can outperform LRs but still trail approaches like delta learning. The later is a negative result showing how fastprop can fail on especially difficult, atypical targets.
All of these fastprop benchmarks are reproducible, and complete instructions for installation, data retrieval and preparation, and training are publicly available on the fastprop GitHub repository at github.com/jacksonburns/fastprop.
Benchmark Methods
The method for splitting data into training, validation, and testing sets varies on a per-study basis and is described in each sub-section. Sampling is performed using the astartes package [28] which implements a variety of sampling algorithms and is highly reproducible. For datasets containing missing target values or invalid SMILES strings, those entries were dropped.
Results for fastprop are reported as the average value of a metric and its standard deviation across a number of repetitions (repeated re-sampling of the dataset). The number of repetitions is chosen to either match referenced literature studies or else increased from two until the performance no longer meaningfully changes. Note that this is not the same as cross-validation. Each section also includes the performance of a zero-layer (i.e. linear regression) network as a baseline to demonstrate the importance of non-linearity in a deep NN.
For performance metrics retrieved from literature it is assumed that the authors optimized their respective models to achieve the best possible results; therefore, fastprop metrics are reported after model optimization using the fastprop train ... --optimize option. When results are generated for this study using Chemprop, the default settings are used except that the number of epochs is increased to allow the model to converge and batch size is increased to match dataset size and speed up training. Chemprop was chosen as a point of comparison throughout this study since it is among the most accessible and well-maintained software packages for molecular machine learning.
When reported, execution time is as given by the unix time command using Chemprop version 1.6.1 on Python 3.8 and includes the complete invocation of Chemprop, i.e. time chemprop_train .... The insignificant time spent manually collating Chemprop results (Chemprop does not natively support repetitions) is excluded. fastprop is run on version 1.0.6 using Python 3.11 and timing values are reported according to its internal time measurement which was verified to be nearly identical to the Unix time command. The coarse comparison of the two packages is intended to emphasize the scaling of LRs and Deep-QSPR and that fastprop is, generally speaking, much faster. All models trained for this study were run on a Dell Precision series laptop with an NVIDIA Quadro RTX 4000 GPU and Intel Xeon E-2286M CPU.
Performance Metrics
The evaluation metrics used in each of these benchmarks are chosen to match literature precedent, particularly as established by MoleculeNet [25], where available. It is common to use scale-dependent metrics that require readers to understand the relative magnitude of the target variables. The authors prefer more readily interpretable metrics such as (Weighted) Mean Absolute Percentage Error (W/MAPE) and are included where relevant.
All metrics are defined according to their typical formulae which are readily available online and are implemented in common software packages. Those presented here are summarized below, first for regression:
Mean Absolute Error (MAE): Absolute difference between predictions and ground truth averaged across dataset; scale-dependent.
Root Mean Squared Error (RMSE): Absolute differences squared and then averaged; scale-dependent.
Mean Absolute Percentage Error (MAPE): MAE except that differences are relative (i.e. divided by the ground truth); scale-independent, range 0 (best) and up.
Weighted Mean Absolute Percentage Error (WMAPE): MAPE except the average is a weighted average, where the weight is the magnitude of the ground truth; scale-independent, range 0 (best) and up.
Coefficient of Determination (R2): Proportion of variance explained; scale-independent, range 0 (worst) to 1 (best).
and classification: - Area Under the Receiver Operating Curve (AUROC, AUC, or ROC-AUC): Summary statistic combining all possible classification errors; scale-independent, range 0.5 (worst, random guessing) to 1.0 (perfect classifier). - Accuracy: Fraction of correct classifications, expressed as either a percentage or a number; scale-independent, range 0 (worst) to 1 (perfect classifier).
Benchmark Results
See Table LABEL:results_table for a summary of all the results. Subsequent sections explore each in greater detail.
| Benchmark | Samples (k) | Metric |
SOTA |
fastprop |
Chemprop |
p |
| QM9 | ~134 | MAE | 0.0047a | 0.0060 | 0.0081a | ~ |
| Pgp | ~1.3 | AUROC | 0.94b | 0.90 | 0.89b | ~ |
| ARA | ~0.8 | Accuracy | 91c | 88 | 82* | 0.083 |
| Flash | ~0.6 | RMSE | 13.2d | 13.0 | 21.2* | 0.021 |
| YSI | ~0.4 | MAE | 22.3e | 25.0 | 28.9* | 0.29 |
| PAH | ~0.06 | R2 | 0.96f | 0.97 | 0.59* | 0.0012 |
a [10] b [29] c [30] d [31] e [32] f [4] * These reference results were generated for this study.
Statistical comparisons of fastprop to Chemprop (shown in the p column) are performed using the non-parametric Wilcoxon-Mann-Whitney Test as implemented in GNumeric. Values are only shown for results generated in this study which are known to be performed using the same methods. Only the results for Flash and PAH are statistically significant at 95% confidence (p<0.05).
QM9
Originally described in Scientific Data [27] and perhaps the most established property prediction benchmark, Quantum Machine 9 (QM9) provides quantum mechanics derived descriptors for many small molecules containing one to nine heavy atoms, totaling ~134k. The data was retrieved from MoleculeNet [25] in a readily usable format. As a point of comparison, performance metrics are retrieved from the paper presenting the UniMol architecture [10] previously mentioned. In that study they trained on only three especially difficult QOIs (homo, lumo, and gap) using scaffold-based splitting (a more challenging alternative to random splitting), reporting mean and standard deviation across 3 repetitions.
fastprop achieves 0.0060 0.0002 mean absolute error, whereas Chemprop achieves 0.00814 0.00001 and the UniMol framework manages 0.00467 0.00004. This places the fastprop framework ahead of previous learned representation approaches but still trailing UniMol. This is not completely unexpected since UniMol encodes 3D information from the dataset whereas Chemprop and fastprop use only 2D. Future work could evaluate the use of 3D-based descriptors to improve fastprop performance in the same manner that UniMol has with LRs. All methods are better than a purely linear model trained on the molecular descriptors, which manages only a 0.0095 0.0006 MAE.
Pgp
First reported in 2011 by Broccatelli and coworkers [33], this dataset has since become a standard benchmark and is included in the Therapeutic Data Commons (TDC) [34] model benchmarking suite. The dataset maps approximately 1.2k small molecule drugs to a binary label indicating if they inhibit P-glycoprotein (Pgp). TDC serves this data through a Python package, but due to installation issues the data was retrieved from the original study instead. The recommended splitting approach is a 70/10/20 scaffold-based split which is done here with 4 replicates.
The model in the original study uses a molecular interaction field but has since been surpassed by other models. According to TDC the current leader [29] on this benchmark has achieved an AUROC of 0.938 0.002 333See the TDC Pgp leaderboard.. On the same leaderboard Chemprop [7] achieves 0.886 0.016 with the inclusion of additional molecular features. fastprop yet again approaches the performance of the leading methods and outperforms Chemprop, here with an AUROC of 0.903 0.033 and an accuracy of 83.6 4.6%. Remarkably, the linear QSPR model outperforms both Chemprop and fastprop, approaching the performance of the current leader with an AUROC of 0.917 0.016 and an accuracy of 83.8 3.9%.
ARA
Compiled by Schaduangrat et al. in 2023 [30], this dataset maps ~0.8k small molecules to a binary label indicating if the molecule is an Androgen Receptor Antagonist (ARA). The reference study introduced DeepAR, a highly complex modeling approach, which achieved an accuracy of 0.911 and an AUROC of 0.945.
For this study an 80/10/10 random splitting is repeated four times on the dataset since no analogous split to the reference study can be determined. Chemprop takes 16 minutes and 55 seconds to run on this dataset and achieves only 0.824 0.020 accuracy and 0.898 0.022 AUROC. fastprop takes only 1 minute and 54 seconds (1 minute and 39 seconds for descriptor calculation) and is competitive with the reference study in performance, achieving a 88.2 3.7% accuracy and 0.935 0.034 AUROC. The purely linear QSPR model falls far behind these methods with a 71.8 6.6% accuracy and 0.824 0.052 AUROC.
Flash
First assembled and fitted to by Saldana and coauthors [31] the dataset (Flash) includes around 0.6k entries, primarily alkanes and some oxygen-containing compounds, and their literature-reported flash point. The reference study reports the performance on only one repetition, but manually confirms that the distribution of points in the three splits follows the parent dataset. The split itself was a 70/20/10 random split, which is repeated four times for this study.
Using a complex multi-model ensembling method, the reference study achieved an RMSE of 13.2, an MAE of 8.4, and an MAPE of 2.5%. fastprop matches this performance, achieving 13.0 2.0 RMSE, 9.0 0.5 MAE, and 2.7% 0.1% MAPE. Chemprop, however, struggles to match the accuracy of either method - it manages an RMSE of 21.2 2.2, an MAE of 13.8 2.1, and a MAPE of 3.99 0.36%. This is worse than the performance of the linear QSPR model, with an RMSE of 16.1 4.0, an MAE of 11.3 2.9, and an MAPE of 3.36 0.77%.
fastprop dramatically outperforms both methods in terms of training time. The reference model required significant manual intervention to create a model ensemble, so no single training time can be fairly identified. fastprop arrived at the indicated performance without any manual intervention in only 30 seconds, 13 of which were spent calculating descriptors. Chemprop, in addition to not reaching the same level of accuracy, took 5 minutes and 44 seconds to do so - more than ten times the execution time of fastprop.
YSI
Assembled by Das and coauthors [32] from a collection of other smaller datasets, this dataset maps ~0.4k molecular structures to a unified-scale Yield Sooting Index (YSI), a molecular property of interest to the combustion community. The reference study performs leave-one-out cross validation to fit a per-fragment contribution model, effectively a training size of >99%, without a holdout set. Though this is not standard practice and can lead to overly optimistic reported performance, the results will be carried forward regardless. The original study did not report overall performance metrics, so they have been re-calculated for this study using the predictions made by the reference model as provided on GitHub 444Predictions are available at this permalink to the CSV file on GitHub.. For comparison fastprop and Chemprop use a more typical 60/20/20 random split and 8 repetitions. Results are summarized in Table LABEL:ysi_results_table.
| Model | MAE | RMSE | WMAPE |
| Reference | 22.3 | 50 | 14 |
| fastprop | 25.0 5.2 | 52 20 | 13.6 1.3 |
| Chemprop | 28.9 6.5 | 63 14 | ~ |
| Linear | 82 39 | 180 120 | 0.47 0.23 |
fastprop again outperforms Chemprop, in this case approaching the overly-optimistic performance of the reference model. Taking into account that reference model has been trained on a significantly larger amount of data, this performance is admirable. Also notable is the difference in training times. Chemprop takes 7 minutes and 2 seconds while fastprop completes in only 42 seconds, again a factor of ten faster. The linear QSPR model fails entirely, performing dramatically worse than all other models.
PAH
Originally compiled by Arockiaraj et al. [4] the Polycyclic Aromatic Hydrocarbons (PAH) dataset contains water/octanol partition coefficients (logP) for 55 polycyclic aromatic hydrocarbons ranging in size from napthalene to circumcoronene. This size of this benchmark is an ideal case study for the application of fastprop. Using expert insight the reference study designed a novel set of molecular descriptors that show a strong correlation to logP, with correlation coefficients ranging from 0.96 to 0.99 among the various new descriptors.
For comparison, fastprop and Chemprop are trained using 8 repetitions of a typical 80/10/10 random split - only 44 molecules in the training data. fastprop matches the performance of the bespoke descriptors with a correlation coefficient of 0.972 0.025. This corresponds to an MAE of 0.19 0.10 and an MAPE of 2.5 1.5%. Chemprop effectively fails on this dataset, achieving a correlation coefficient of only 0.59 0.24, an MAE of 1.04 0.33 (one anti-correlated outlier replicate removed). This is worse even than the purely linear QSPR model, which manages a correlation coefficient of 0.78 0.22, an MAE of 0.59 0.22, and an RMSE of 0.75 0.32. Despite the large parameter size of the fastprop model relative to the training data, it readily outperforms Chemprop in the small-data limit.
For this unique dataset, execution time trends are inverted. fastprop takes 1 minute and 43 seconds, of which 1 minute and 31 seconds were spent calculating descriptors for these unusually large molecules. Chemprop completes in 1 minute and 16 seconds, faster overall but much slower compared with the training time of fastprop without descriptor calculation.
Limitations and Future Work
Negative Results
Delta Learning with Fubrain
First described by Esaki and coauthors, the Fraction of Unbound Drug in the Brain (Fubrain) dataset is a collection of about 0.3k small molecule drugs and their corresponding experimentally measured unbound fraction in the brain, a critical metric for drug development [23]. This specific target in combination with the small dataset size makes this benchmark highly relevant for typical QSPR studies, particular via delta learning. DeepDelta [15] performed a 90/0/10 cross-validation study of the Fubrain dataset in which the training and testing molecules were intra-combined to generate all possible pairs and then the differences in the property 555Although the original Fubrain study reported untransformed fractions, the DeepDelta authors confirmed via GitHub that DeepDelta was trained on log base-10 transformed fraction values, which is replicated here. were predicted, rather than the absolute values, increasing the amount of training data by a factor of 300.
DeepDelta reported an RMSE of 0.830 0.023 at predicting differences, whereas a typical Chemprop model trained to directly predict property values was only able to reach an accuracy of 0.965 0.019 when evaluated on its capacity to predict property differences. fastprop is able to outperform Chemprop, though not DeepDelta, achieving an RMSE of 0.930 0.029 when using the same splitting procedure above. It is evident that delta learning is still a powerful technique for regressing small datasets.
For completeness, the performance of Chemprop and fastprop on Fubrain are also compared to the original study. The study that first generated this dataset used mordred descriptors but as is convention they strictly applied linear modeling methods. Using both cross validation and and external test sets, they had an effective training/validation/testing split of 0.64/0.07/0.28 which will be repeated 4 times here for comparison. All told, their model achieved an RMSE of 0.53 averaged across all testing data. In only 39 seconds, of which 31 are spent calculating descriptors, fastprop far exceeds the reference model with an RMSE of 0.207 0.024. This also surpasses Chemprop, itself outperforming the reference model with an RMSE of 0.223 0.036.
fastprop Fails on QuantumScents
Compiled by Burns and Rogers [35], this dataset contains approximately 3.5k SMILES and 3D structures for a collection of molecules labeled with their scents. Each molecule can have any number of reported scents from a possible 113 different labels, making this benchmark a a Quantitative Structure-Odor Relationship. Due to the highly sparse nature of the scent labels a unique sampling algorithm (Szymanski sampling [36]) was used in the reference study and the exact splits are replicated here for a fair comparison.
In the reference study, Chemprop achieved an AUROC of 0.85 with modest hyperparameter optimization and an improved AUROC of 0.88 by incorporating the atomic descriptors calculated as part of QuantumScents. fastprop is incapable of incorporating atomic features, so they are not included. Using only the 2D structural information, fastprop falls far behind the reference study with an AUROC of only 0.651 0.005. Even when using the high-quality 3D structures and calculating additional descriptors (demonstrated in the GitHub repository), the performance does not improve.
The exact reason for this failure is unknown. Possible reasons include that the descriptors in mordred are simply not correlated with this target, and thus the model struggles to make predictions. This is a fundamental drawback of this fixed representation method - whereas a LR could adapt to this unique target, fastprop fails.
Execution Time
fastprop is consistently faster to train than Chemprop when using a GPU, helping exploit the ‘time value’ of data. Note that due to the large size of the FNN in fastprop it can be slower than small Chemprop models when training on a CPU since Chemprop uses a much smaller FNN and associated components.
There is a clear performance improvement to be had by reducing the number of descriptors to a subset of only the most important. Future work can address this possibility to decrease time requirements for both training by reducing network size and inference by decreasing the number of descriptors to be calculated for new molecules. This has not been done in this study for two reasons: (1) to emphasize the capacity of the DL framework to effectively perform feature selection on its own via the training process, de-emphasizing unimportant descriptors; (2) as discussed above, training time is small compared to dataset generation time, or even compared to to the time it takes to compute the descriptors using mordred.
Coverage of Descriptors
fastprop is fundamentally limited by the types of chemicals which can be uniquely described by the mordred package. Domain-specific additions which are not just derived from the descriptors already implemented will be required to expand its application to new domains.
For example, in its current state mordred does not include any connectivity based-descriptors that reflect the presence or absence of stereocenters. While some of the 3D descriptors it implements could implicitly reflect sterochemistry, more explicit descriptors like the Stereo Signature Molecular Descriptor [37] may prove helpful in the future if re-implemented in mordred.
Interpretability
Though not discussed here for the sake of length, fastprop already contains the functionality to perform feature importance studies on trained models. By using SHAP values [38] to assign a scalar ‘importance’ to each of the input features, users can determine which of the mordred descriptors has the largest impact on model predictions. The utility of these values can be explored in greater detail on a case-by-case basis.
Availability
Project name: fastprop
Project home page: github.com/jacksonburns/fastprop
Operating system(s): Platform independent
Programming language: Python
Other requirements: pyyaml, lightning, mordredcommunity, astartes
Declarations
Availability of data and materials
fastprop is Free and Open Source Software; anyone may view, modify, and execute it according to the terms of the MIT license. See github.com/jacksonburns/fastprop for more information.
All data used in the Benchmarks shown above is publicly available under a permissive license. See the benchmarks directory at the fastprop GitHub page for instructions on retrieving each dataset and preparing it for use with fastprop, where applicable.
Competing interests
None.
Funding
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0023112.
Authors’ contributions
Initial ideation of fastprop was a joint effort of Burns and Green. Implementation, benchmarking, and writing were done by Burns.
Acknowledgements
The authors acknowledge Haoyang Wu, Hao-Wei Pang, and Xiaorui Dong for their insightful conversations when initially forming the central ideas of fastprop.
Disclaimer
This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
Cited Works
re 1. Muratov EN, Bajorath J, Sheridan RP, et al (2020) QSAR without borders. Chem Soc Rev 49:3525–3564. https://doi.org/10.1039/D0CS00098A
pre 2. Wiener H (1947) Structural determination of paraffin boiling points. Journal of the American Chemical Society 69:17–20. https://doi.org/10.1021/ja01193a005
pre 3. Estrada’ E, Torres’ L, Rodriguez’ L (1998) An atom-bond connectivity index: Modelling the.enthalpy of formation of alkanes. Indian Journal of Chemistry 37:849–855
pre 4. Arockiaraj M, Paul D, Clement J, et al (2023) Novel molecular hybrid geometric-harmonic-zagreb degree based descriptors and their efficacy in QSPR studies of polycyclic aromatic hydrocarbons. SAR and QSAR in Environmental Research 34:569–589. https://doi.org/10.1080/1062936x.2023.2239149
pre 5. Ma J, Sheridan RP, Liaw A, et al (2015) Deep neural nets as a method for quantitative structure-activity relationships. Journal of Chemical Information and Modeling 55:263–274. https://doi.org/10.1021/ci500747n
pre 6. Coley CW, Barzilay R, Jaakkola TS, et al (2017) Prediction of organic reaction outcomes using machine learning. ACS Central Science 3:434–443. https://doi.org/10.1021/acscentsci.7b00064
pre 7. Yang K, Swanson K, Jin W, et al (2019) Analyzing learned molecular representations for property prediction. Journal of Chemical Information and Modeling 59:3370–3388. https://doi.org/10.1021/acs.jcim.9b00237
pre 8. Heid E, Greenman KP, Chung Y, et al (2024) Chemprop: A machine learning package for chemical property prediction. Journal of Chemical Information and Modeling 64:9–17. https://doi.org/10.1021/acs.jcim.3c01250
pre 9. Song Y, Zheng S, Niu Z, et al (2021) Communicative representation learning on attributed molecular graphs. In: Proceedings of the twenty-ninth international joint conference on artificial intelligence. Yokohama, Yokohama, Japan
pre 10. Zhou G, Gao Z, Ding Q, et al (2023) Uni-mol: A universal 3D molecular representation learning framework. In: The eleventh international conference on learning representations
pre 11. Chen J, Schwaller P (2023) Molecular hypergraph neural networks. https://doi.org/10.48550/arXiv.2312.13136
pre 12. Zhao B, Xu W, Guan J, Zhou S (2023) Molecular property prediction based on graph structure learning. arXiv preprint arXiv:231216855
pre 13. Wang Z, Jiang T, Wang J, Xuan Q (2024) Multi-modal representation learning for molecular property prediction: Sequence, graph, geometry. https://doi.org/10.48550/ARXIV.2401.03369
pre 14. Zhu Y, Chen D, Du Y, et al (2024) Molecular contrastive pretraining with collaborative featurizations. Journal of Chemical Information and Modeling 64:1112–1122. https://doi.org/10.1021/acs.jcim.3c01468
pre 15. Schaduangrat N, Anuwongcharoen N, Charoenkwan P, Shoombuatong W (2023) DeepAR: A novel deep learning-based hybrid framework for the interpretable prediction of androgen receptor antagonists. Journal of Cheminformatics 15:50. https://doi.org/10.1186/s13321-023-00721-z
pre 16. Tilborg D van, Brinkmann H, Criscuolo E, et al (2024) Deep learning for low-data drug discovery: Hurdles and opportunities. ChemRxiv. https://doi.org/10.26434/chemrxiv-2024-w0wvl
pre 17. Moriwaki H, Tian Y-S, Kawashita N, Takagi T (2018) Mordred: A molecular descriptor calculator. Journal of Cheminformatics 10:4. https://doi.org/10.1186/s13321-018-0258-y
pre 18. McGibbon M, Shave S, Dong J, et al (2023) From intuition to AI: evolution of small molecule representations in drug discovery. Briefings in Bioinformatics 25:bbad422. https://doi.org/10.1093/bib/bbad422
pre 19. Falcon W, The PyTorch Lightning team (2019) PyTorch Lightning
pre 20. Weininger D (1988) SMILES, a chemical language and information system. 1. Introduction to methodology and encoding rules. Journal of Chemical Information and Computer Sciences 28:31–36. https://doi.org/10.1021/ci00057a005
pre 21. Comesana AE, Huntington TT, Scown CD, et al (2022) A systematic method for selecting molecular descriptors as features when training models for predicting physiochemical properties. Fuel 321:123836. https://doi.org/https://doi.org/10.1016/j.fuel.2022.123836
pre 22. Wu J, Wang S, Zhou L, et al (2020) Deep-learning architecture in QSPR modeling for the prediction of energy conversion efficiency of solar cells. Industrial and Engineering Chemistry Research 59:18991–19000. https://doi.org/10.1021/acs.iecr.0c03880
pre 23. Esaki T, Ohashi R, Watanabe R, et al (2019) Computational model to predict the fraction of unbound drug in the brain. Journal of Chemical Information and Modeling 59:3251–3261. https://doi.org/10.1021/acs.jcim.9b00180
pre 24. Yalamanchi KK, Kommalapati S, Pal P, et al (2023) Uncertainty quantification of a deep learning fuel property prediction model. Applications in Energy and Combustion Science 16:100211. https://doi.org/https://doi.org/10.1016/j.jaecs.2023.100211
pre 25. Wu Z, Ramsundar B, Feinberg EN, et al (2018) MoleculeNet: A benchmark for molecular machine learning. https://doi.org/10.48550/arXiv.1703.00564
pre 26. Ramakrishnan R, Hartmann M, Tapavicza E, von Liliendfeld O (2015) Electronic spectra from TDDFT and machine learning in chemical space. The Journal of Chemical Physics 143: https://doi.org/10.1063/1.4928757
pre 27. Ramakrishnan R, Dral P, Rupp M, von Liliendfeld O (2014) Quantum chemistry structures and properties of 134 kilo molecules. Scientific Data 1: https://doi.org/10.1038/sdata.2014.22
pre 28. Burns J, Spiekermann K, Bhattacharjee H, et al (2023) Machine learning validation via rational dataset sampling with astartes. Journal of Open Source Software 8:5996. https://doi.org/10.21105/joss.05996
pre 29. Notwell JH, Wood MW (2023) ADMET property prediction through combinations of molecular fingerprints. https://doi.org/10.48550/arXiv.2310.00174
pre 30. Schaduangrat N, Anuwongcharoen N, Charoenkwan P, Shoombuatong W (2023) DeepAR: A novel deep learning-based hybrid framework for the interpretable prediction of androgen receptor antagonists. Journal of Cheminformatics 15:50. https://doi.org/10.1186/s13321-023-00721-z
pre 31. Saldana DA, Starck L, Mougin P, et al (2011) Flash point and cetane number predictions for fuel compounds using quantitative structure property relationship (QSPR) methods. Energy & Fuels 25:3900–3908. https://doi.org/10.1021/ef200795j
pre 32. Das DD, St. John PC, McEnally CS, et al (2018) Measuring and predicting sooting tendencies of oxygenates, alkanes, alkenes, cycloalkanes, and aromatics on a unified scale. Combustion and Flame 190:349–364. https://doi.org/10.1016/j.combustflame.2017.12.005
pre 33. Broccatelli F, Carosati E, Neri A, et al (2011) A novel approach for predicting p-glycoprotein (ABCB1) inhibition using molecular interaction fields. Journal of Medicinal Chemistry 54:1740–1751. https://doi.org/10.1021/jm101421d
pre 34. Huang K, Fu T, Gao W, et al (2021) Therapeutics data commons: Machine learning datasets and tasks for therapeutics. CoRR abs/2102.09548: https://doi.org/10.48550/arXiv.2102.09548
pre 35. Burns JW, Rogers DM (2023) QuantumScents: Quantum-mechanical properties for 3.5k olfactory molecules. Journal of Chemical Information and Modeling 63:7330–7337. https://doi.org/10.1021/acs.jcim.3c01338
pre 36. Szymański P, Kajdanowicz T (2017) A network perspective on stratification of multi-label data. In: Luís Torgo PB, Moniz N (eds) Proceedings of the first international workshop on learning with imbalanced domains: Theory and applications. PMLR, pp 22–35
pre 37. Carbonell P, Carlsson L, Faulon J-L (2013) Stereo signature molecular descriptor. Journal of Chemical Information and Modeling 53:887–897. https://doi.org/10.1021/ci300584r
pre 38. Lundberg S, Lee S-I (2017) A unified approach to interpreting model predictions. https://doi.org/10.48550/ARXIV.1705.07874
p